OpenStudy (anonymous):

What are the x-intercept(s) of the graph of y + 6 = x2 − x? a. (−2, 0) and (3, 0) b.(−2, 0) and (−3, 0) c.(2, 0) and (−3, 0) d.(2, 0) and (3, 0)

jimthompson5910 (jim_thompson5910):

Plug in y = 0 and solve for x y + 6 = x^2 − x 0 + 6 = x^2 − x 6 = x^2 - x x^2 - x - 6 = 0 (x-3)(x+2) = 0 See where to go from here?